OpenStudy (anonymous):

Write the slope intercept equation for the line thant passes through (-7,-1) and has slope -6/7. Please show work. YOU WILL GET THE MEDAL

OpenStudy (anonymous):

\[y = mx+c\] Is the Slope-Intercept form !

OpenStudy (anonymous):

\[m= -\frac{6}{7}\]

OpenStudy (anonymous):

use this form first: y-y1=m(x-x1)

OpenStudy (anonymous):

Then Just Plug in Points an get C

OpenStudy (anonymous):

you got me lost

OpenStudy (anonymous):

\[C = -7\]

OpenStudy (anonymous):

\[Hence\:\:\:\:\:\:\:\: y = -\frac{6}{7}x -7\]

OpenStudy (anonymous):

ok where do the -1 come in that

OpenStudy (amistre64):

the expanded version of the slope intercept form is: y = mx -mPx +Py; where m=slope, P(x,y) = Px, Py

OpenStudy (anonymous):

that is makeing me more confused

OpenStudy (amistre64):

i dont see how it confuses you. It is a simplified formula that removes a few steps ...